OpenStudy (anonymous):

Simplify. Write you answer using only POSITIVE exponents. SHOW ALL WORK FOR THIS PROBLEM (-5x^2* y^-2*z^5/25x^9*yz^-1)^-2 Please help me my answer z^6/5x^7y^3 Simplify. Write you answer using only POSITIVE exponents. SHOW ALL WORK FOR THIS PROBLEM (-5x^2* y^-2*z^5/25x^9*yz^-1)^-2 Please help me my answer z^6/5x^7y^3 @Mathematics

OpenStudy (anonymous):

\[({-5x^2y^{-2}z^5}/{25x^9yz^{-1}})^{-2}=(-z^6/5x^7y^3)^{-2}=(5x^7y^3/-z^6)^2=(25x^{14}y^6/z^{12})\]